OpenStudy (mackenzie2013):

A function has the function rule y = -7x - 2. If the input is -4, what is the output? 26 -30 -26 14

OpenStudy (anonymous):

y is the output and x is the input. plug in -4 for x in the function rule to get the output.

OpenStudy (mackenzie2013):

Thank you!

OpenStudy (bookworm14):

substitute -4 for x y=-7(-4) -2 or -7(-4)-2 = y -7*-4=28 ..... 28-2 =26 26 is your answer Hope this helps :) Bree